OpenStudy (anonymous):

a farmer said to another farmer , if you give me one of your horses, we'll each have the same number. the other replied, if you give me one horse, i will have twice as many as you. How many horses did each have to begin with? I know it is 5 and 7, but i do not know how to solve it algebraically

OpenStudy (nikato):

Ok, let a be farmer 1 and b be 2. U can write some equations with the given info. a+1=b-1 b+1=2a

OpenStudy (nikato):

For the first equation, add 1 to both sides so a+2=b

OpenStudy (nikato):

Wait second equation is b+1=2(a-1)

OpenStudy (nikato):

Substitute this^ to the second equation a+2+1=2(a-1) Simplify a+3=2a-2 and then a=5

OpenStudy (nikato):

Then substitute a back to one of the original to get b

OpenStudy (nikato):

Do u have any questions about this?^ @jesslovespink96

OpenStudy (anonymous):

so i substitute the second equation with the one that says simplify under it?

OpenStudy (nikato):

No, simply is just me simplifying. U see my second post and the equation after add 1 to both sides, the one with b=a+2. U want substitr the value of b into the second original equation

OpenStudy (nikato):

And after substitution u should get the equation above simplify

OpenStudy (anonymous):

oh yes i got the right answer, thank you very much!

OpenStudy (nikato):

Yea, no problem

OpenStudy (whpalmer4):

Here's another version: \[a + 1 = b - 1\]\[2(a-1) = b+1\] Those are the two statements made between the farmers, translated to equations as before. Now we will expand the second equation and multiply each term by -1: \[a + 1 = b - 1\]\[-2a+2 = -b - 1\]Now I add the two equations together, down the columns: \[a - 2a + 1 + 2 = b - b - 1 - 1\]\[-a + 3 = -2\]\[a = 5\]Substitute our value of \(a\) back into either of the original equations and solve for \(b\): \[5+1=b-1\]\[6=b-1\]\[7=b\]or\[2(5-1) = b+1\]\[8=b+1\]\[7=b\]

OpenStudy (whpalmer4):

I multiplied by -1 so that when I added the equations together, the \(b\) terms cancelled each other out, leaving me an equation just in terms of \(a\) to solve.